OpenStudy (anonymous):

pls help me answer dis ryt now. it says on the problem.. solve s=-1/2gt(squared) + vt for g pls help me ryt now.. sorry i dnt how to put exponent 2..so i just put squared..

OpenStudy (anonymous):

Do you mean: \[s = -\frac{1}{2}g(t^2) + vt\]

OpenStudy (anonymous):

If you do then: \[2s - 2vt = -g(t^2)\] \[g = \frac{2(vt-s)}{t^2}\]

OpenStudy (anonymous):

uhh @mathsboy , no parenthesis outside the t squared

OpenStudy (nilankshi):

???????????

OpenStudy (anonymous):

I used the parentheses to make sure you meant you're not squaring the whole of (gt)

OpenStudy (nilankshi):

lol

OpenStudy (anonymous):

ohh olryt . so can you help me solve it? plsss

OpenStudy (anonymous):

I've done as much as I can, unless you have values of s, t and v it can't be taken any further.

OpenStudy (anonymous):

and ya its my pic to those hu ar asking

OpenStudy (anonymous):

The final answer is \[g = \frac{2(vt-s)}{t^2}\]

OpenStudy (nilankshi):

i was asking it

OpenStudy (anonymous):

ohh olryt. SOO THANK YOU... your soo smart ..

OpenStudy (nilankshi):

u ar in which grade

OpenStudy (nilankshi):

???????????

OpenStudy (anonymous):

@ mathsboy, ar yu online evryday?

OpenStudy (anonymous):

can i ask another math question again?